OpenStudy (anonymous):

Write 56 1/4%as a fraction in simpliest form

OpenStudy (imnotgoodatmath):

I Think 9/16

OpenStudy (alekos):

i agree

OpenStudy (texaschic101):

56 1/4% = 56.25% = 5625/10,000 = 9/16 yep....me too